OpenStudy (idec):

I am really bad at justifying steps within algebra and I need help with this problem.

OpenStudy (idec):

Here is a screenshot of the question.

OpenStudy (osprey):

multiply out the bracket

OpenStudy (osprey):

rearrange terms on the left hand side

OpenStudy (osprey):

group 4+3

OpenStudy (osprey):

add 4+3=7

OpenStudy (idec):

I need to name the property of the step used, such as if it were the distributive or commutative property.

OpenStudy (osprey):

i half thought that was coming. ok distributive is sort of multiplying out (distributing things is giiving them out). commutative is when you can reverse things so 2x3 is the same as 3x2

OpenStudy (unklerhaukus):

A + B = B + A (commutative property of addition)

OpenStudy (unklerhaukus):

A + (B + C) = (A + B) + C (associative property of addition)

OpenStudy (unklerhaukus):

n(A + B) = nA + nB (distributive property of multiplication)